Presentation is loading. Please wait.

Presentation is loading. Please wait.

Focus Coil Update Roy Preece / Tom Bradshaw / Mike Courthold STFC – RAL VC 17 th Jan 2013.

Similar presentations

Presentation on theme: "Focus Coil Update Roy Preece / Tom Bradshaw / Mike Courthold STFC – RAL VC 17 th Jan 2013."— Presentation transcript:

1 Focus Coil Update Roy Preece / Tom Bradshaw / Mike Courthold STFC – RAL VC 17 th Jan 2013

2 Since last meeting Gas input from bottle pack – 1.05 bar Level sensor operating but not able to communicate with the control system. System cooling over the Christmas period By the New year the level was reported as being 100%. Pressure feedback heater systems operating with a regulation of 0.005 bar. Level reduced to 80% using the pressure regulation heaters and the exhaust poppet valve. Pressure regulated at 1.14 bar with a gas bottle backup set to 1.05 bar. Good news : Easily demonstrated zero boil off conditions Good news : Control system and monitoring works well

3 Review and Issues Over the holiday period there were issues surrounding the remote monitoring of the system This was complicated by the loss of air conditioning in a room somewhere …. – this took down several of the PPD computers Other software issues concerned versions of the software running on different computers Although the data was archived it was not always accessible through the archiver. Where are the computers that we are using located ? What networks are they on? Who is responsible for this kit ? Over the holiday period it was difficult to resolve these issues

4 Review and Issues A pre powering review was held on the 8 th Jan. – Risk and method statements nominally agreed (some amendments) – Pressure relief mainly reliant on the bursting disc. Relief valve with higher flow rate added to the system. – Ready to go forward with final preparations and powering. Issues – Current lead heaters not working Found to have very high resistance to new installed. ON / OFF reporting on control screen sorted. – Connection to ground found on voltage tap Problem rectified and full connectivity check carried out.

5 Review and Issues Powering of the coils started on 11 th Jan – Quench detection system tripping due to “Step function” from the power supply controller, AMI 430. – Maximum current output set to 5A and QD system taken out of the control loop. – Current of 5A in the coils but, voltage spikes evident while the controller is trying to control the current. 0.14T seen at the centre of the magnet bore. – The cause of the “Step function” is unknown so AMI are contacted for assistance. – Testing operations were stopped at this point and the magnet was left in a safe but cold state.

6 Review and Issues DL engineers on site at RAL working on the controller problem today. (More up to date information in the meeting) – Earthing of the controller checked as per AMI suggestion with no change. – Further investigations into addition earth loops – Change the controller for a spare unit to ascertain if the same issue is seen. – Blah Blah Blah Blah………. – None of this was needed because the problem was found to be the instructions from AMI are incorrect and at the wrong end of the scale. 95% stability setting required NOT 20%.

7 Plans Ramp the magnet to 5A checking the voltage and QD operation Ramp to 50A and hold. – Take thermal measurements to check the thermal model – Take magnetic field measurements to check the field analysis Induce a quench in the magnet – This will check all QD, dump circuitry, logging etc – Full pressure system checks Attach the gas bottle pack to the input – Re-cool the system – Condense He into cooling channels LHe delivery week of 21 st Jan (next week) – Refill and stabilise – Continue testing to full current in Solenoid mode – 114A – Thermal and magnetic measurements – Re-configure for Flip mode – Power to full current, with stage points, to 225A

Download ppt "Focus Coil Update Roy Preece / Tom Bradshaw / Mike Courthold STFC – RAL VC 17 th Jan 2013."

Similar presentations

Ads by Google